OpenStudy (anonymous):

can some one help me whit slope my points are (-2,2), and (2,-2).

OpenStudy (anonymous):

it wants me to arrange them y=mx+b

OpenStudy (imstuck):

THis is the slope formula:\[m=\frac{ y _{2}-y _{1} }{x _{2}-x _{1} }\]Filling in your points you will get this:\[m=\frac{ -2-2 }{ 2-(-2) }\]Can you do that math?

OpenStudy (imstuck):

AFter you find the slope you can write it in that form, which is slope-intercept form.

OpenStudy (imstuck):

Can you find the slope from what I gave you?

OpenStudy (anonymous):

0

OpenStudy (imstuck):

No, that's not right. Here, like this:

OpenStudy (imstuck):

\[\frac{ -2-2 }{ 2-(-2) }=\frac{ -4 }{ 4 }=-1\]

OpenStudy (anonymous):

yeah my bad -1

OpenStudy (imstuck):

Ok, good. Now you have to use that slope and one of the points in the point-slope formula to get the equation of the line. Do you know the point-slope formula?

OpenStudy (anonymous):

y=mx+b

OpenStudy (imstuck):

No, that's slope-intercept. You have to use the point-slope formula to get the equation, then rearrange it to put it into slope-intercept. Do you know the point-slope formula?

OpenStudy (anonymous):

no thats all it shows me

OpenStudy (imstuck):

Ok, first thing to do is pick one of the points they gave you...either the (-2,2) or the (2,-2). I'm going to pick the first one, since they both are just about the same with one having a positive and one having a negative. This is the point-slope formula:\[y-y _{1}=m(x-x _{1})\]where y1 is your y coordinate from the point you chose, and x1 is the x coordinate from the point you chose. I will fill it in and show you how to do it.

OpenStudy (imstuck):

I meant that both the points have a positive and a negative. I worded that badly. Sorry.

OpenStudy (anonymous):

no prob

OpenStudy (imstuck):

\[y-y _{1}=m(x-x _{1})\]\[y-2=-1(x-(-2))\]"m" is your slope of -1 that you found earlier.

OpenStudy (imstuck):

Going on, you get this:

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

\[y-2=-1(x+2)\]\[y-2=-1x-2\]Now here is where you move stuff around to get it into slope-intercept, or y = mx + b, form.

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

y-2+2=-1x-2+2 y=-1x or y = -x

OpenStudy (imstuck):

Because the 2's cancel each other out on both sides, there is no need to put it in, cuz if you did it would be as a 0, which you can do if you want. If so, the equation would look like this: y = -1x + 0

OpenStudy (imstuck):

When you are given 2 points and told to write it in slope-intercept form, first you have to find the slope, then you pick a point and use the point-slope formula, then you rearrange it to put it into slope-intercept form. 3 steps to it each time.